What is f(x + 1) if f(x)= 6x^3 - 3x^2 + 4x - 9
I REALLY appreciate the help!

Answer by stanbon(48558)

What is f(x + 1) if f(x)= 6x^3 - 3x^2 + 4x - 9
f(x+1) = 6(x+1)^3 - 3(x+1)^2 + 4(x+1) - 9
f(x+1) = 6[x^3+3x^2+3x+1] -3[x^2+2x+1] + 4x+4 -9
f(x+1) = 6x^3 +15x^2 +16x -2
==============
Cheers,
Stan H.